OpenStudy (anonymous):

Let f(x) = x^3 - 6x^2 + 8x + 2. For what value(s) of x is f(x) = 2?

OpenStudy (anonymous):

if thats the case let f(x)=2 be f(2) plugging in to where a x's are

OpenStudy (anonymous):

\[x^{3} - 6x^{2} + 8x + 2 = 2\]\[x^{3} - 6x^{2} + 8x = 0\]\[x(x^{2} - 6x + 8) = 0\]\[x(x -4)(x - 2) = 0\] Can you finish from there?

OpenStudy (anonymous):

thanks i got it now

OpenStudy (anonymous):

np